# Optimal control under unknown intensity with Bayesian learning

### v1

@@ Abstract
We consider an optimal control problem inspired by neuroscience, where the dynamics is driven by a Poisson process with a controlled stochastic intensity and an uncertain parameter. Given a prior distribution for the unknown parameter, we describe its evolution according to Bayes’ rule. We reformulate the optimization problem using Girsanov’s theorem and establish a dynamic programming principle. Finally, we characterize the value function as the unique viscosity solution to a finite-dimensional Hamilton-Jacobi-Bellman equation, which can be solved numerically.
@@ 1 Introduction
Let $(Y_{t})$ represent the membrane potential of a spiking neuron. The neuron emits “spikes” randomly at rate $f_{\Lambda}(Y_{t})$. When a spike occurs, the potential is reset to zero. Finally, between the jumps, the dynamics is given by the ODE
@@
In this model, the functions $(\lambda,y)\mapsto f_{\lambda}(y)\in\mathbb{R}_{+}$ and $y\mapsto b(y)\in\mathbb{R}$ are deterministic and known. In addition, $\Lambda$ is a random variable with known distribution $\mu$, supported on $\mathbb{R}_{+}$, but the realization of $\Lambda$ is unknown. Our objective is to estimate as accurately as possible the value of $\Lambda$. In addition, the input current $(\gamma_{t})$ is a control and can be chosen in order to improve the estimation. Crucially, we only observe the jumping times of $(Y_{t})$ and therefore $(\gamma_{t})$ is only allowed to causally depend on the spiking times up to time $t$. The question we answer in this article can be summarized as follows:
How to choose the stochastic control $(\gamma_{t})$ to optimally estimate the unknown parameter?
This leads us to consider an optimal control problem driven by a counting process with unknown intensity. Our approach is applicable beyond neuroscience, where optimal control over processes with unknown parameters is required, such as in online learning scenarios. We present a mathematical framework to tackle this class of questions.
Such problems have been widely studied in the discrete-time stochastic optimal control literature (see, e.g., [6, 7]). A key approach is to assign a prior distribution to the unknown parameter, which is updated at each time step using Bayes’ rule. In the continuous-time setting, particularly within the impulse control framework, these issues have been addressed by [2] in a Brownian framework and extended by [1] in a Poisson framework. As in the discrete-time setting, a Bayesian framework is used: the unknown parameter is sampled from a prior distribution and a stochastic process in the space of probability measures is introduced, known as the posterior distribution. This approach leads to a dynamic programming principle and a characterization of the value function via a quasi-variational parabolic equation interpreted in the sense of viscosity solutions.
In [8] (lecture given at Collège de France) and [4] (with an application in asset management), the authors consider instances of the following diffusive case:
@@
In this model, one observes the process $(X_{t})$ and the parameter $\Lambda$ is unknown. A key idea introduced in these works is to reformulate the problem using Girsanov’s theorem. This permits removing the drift and to make explicit the dependence of the problem on the unknown parameter via the exponential martingale involved in Girsanov’s transformation. In addition, in this linear case in $\Lambda$ or for specific classes of prior distribution known as conjugate families, they derive a finite dimensional HJB equation satisfied by the value function.
@@
In the setting we introduce, a new mathematical challenge arises because the conditional distribution of the unknown parameters evolves continuously and also exhibits jumps: the jumps of the conditional distribution occurs precisely at the jumping times of the underlying point process; with intensity depending on the unknown parameter. Another new important difficulty appear: when the control $(\gamma_{t})$ is modified, the jumps of the process are also modified. As the controller is only allowed to observe the spiking times, it is not clear to know at first sight what are the admissible controls. In particular, it is not possible to use the same control $(\gamma_{t})$ for two different initial conditions $y$ and $y^{\prime}$ of $(Y_{t})$ because changing the initial condition might change the spiking times and therefore a control which was valid for $y$ is not valid anymore for $y^{\prime}$. This feature makes particularly difficult to obtain apriori regularity estimates on the value function. As in [8, 4], Girsanov’s theorem plays an important role in our work in addressing these difficulties. Another difficulty is due to the fact that the measure space is infinite-dimensional, and the posterior distribution evolves continuously over time on this space. In the case where the intensity of the original point process is linearly dependent on the unknown parameter, we show that the posterior distribution is confined to a two-dimensional subspace of $\mathbb{R}^{2}$. We further show that the value function associated with the optimal control problem is the unique solution to a specific finite-dimensional Hamilton-Jacobi-Bellman (HJB) equation in the viscosity sense.
We now describe briefly our main contributions and the organization of this paper.
@@
The control problem is introduced in Section 2.1, where we also characterize the admissible controls and show that the optimization problem is independent of the probability space we started with. In Section 2.2, we reformulate the problem using Girsanov’s theorem for point processes with stochastic intensity and show the equivalence between the original and the new formulation. In Section 2.3, the value function is introduced, and in Section 2.4, we show how to obtain a finite-dimensional representation of the value function provided that function $f_{\lambda}(y)$ is linear in $\lambda$, that is $f_{\lambda}(y)=\lambda g(y)$. This is analogous to the finite-dimensional reduction obtained in [8, 4] in the diffusive setting with a linear drift. Thereafter, we concentrate our study on this finite-dimensional problem.
@@
In Section 3.1, we prove that the value function is Lipschitz continuous with respect to the space variables. To proceed, we study in detail the exponential martingale involved in the Girsanov’s transformation. Using these apriori regularity estimates, we prove in Section 3.2 a first dynamic programming principle with deterministic stopping times. Leveraging this result, we show in Section 3.3 that the value function is continuous with respect to time. Finally, in Section 3.4 we provide a general dynamic programming principle that allows for general stopping times.
@@
In Section 4, we show that the value function is the unique viscosity solution of a certain HJB equation, see equation (4.1). This first order HJB equation has zero order terms which are inherent to the jumps in the underlying dynamics. We rely on the methodology of [9] to handle these terms. A specific difficulty appears in the proof of our comparison result, due to the fact that the controls are not bounded, neither are the space variables.
@@
Finally, in Section 5, we present several numerical illustrations and explain the behavior of the optimal control.
@@ 2.1 The optimization problem
Let $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$ be a probability measure supported on $\mathbb{R}_{+}$ with a finite second moment. Consider functions $b$ and $f$ such that:
@@
The function $b:\mathbb{R}\rightarrow\mathbb{R}$ is Lipschitz and $f:\mathbb{R}_{+}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ is continuous and bounded for each $\lambda\geq 0$:
$\sup_{y\in\mathbb{R}}f_{\lambda}(y)<\infty.$
Let $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t}),\mathbb{P})$ be a filtered probability space satisfying the usual conditions. Fix a time horizon $T>0$. Consider $(\gamma_{t})_{t\in[0,T]}$ an $({\mathcal{F}}_{t})$-progressively measurable process such that
@@
We consider a triple $(\Lambda,(N_{t}),(Y_{t}))$ which is a solution of:
@@
We assume that the probability space is large enough so that for each $\gamma$ as above, there exists at least one triple $(\Lambda,(N_{t}),(Y_{t}))$ defined on this probability space, satisfying the three conditions above.
Let ${\mathcal{F}}^{N}_{t}$ denote the canonical filtration of $(N_{t})_{t\geq 0}$:
${\mathcal{F}}^{N}_{t}=\sigma\{N_{s},~{}s\leq t\}.$
Let $(\tau_{k})_{k\geq 1}$ denote the successive jump times of $(N_{t})_{t\geq 0}$. We set by convention that $\tau_{0}=0$.
Definition 2.3.
A control $\gamma$ is said to be admissible if it is predictable with respect to the filtration $({\mathcal{F}}^{N}_{t})_{t\geq 0}$.
We denote by ${\mathcal{A}}$ the set of admissible controls. Furthermore, we consider the posterior distribution of $\Lambda$ given the observed spiking times, denoted by:
$M^{\gamma}_{t}:={\mathcal{L}}(\Lambda~{}|~{}{\mathcal{F}}^{N}_{t}).$
Let $\kappa>0$. Our optimization problem is formulated as follows:
In other words, our goal is to minimize the variance of the posterior distribution; the quadratic cost in the control is a regularization to prevent for too large controls. We now rewrite this optimization problem in a more explicit way, specifically to demonstrate that the value defined by (2.1) is independent of the probability space $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t}),\mathbb{P})$. The following proposition characterizes the admissible controls.
Proposition 2.4.
Let $\gamma$ be a ${\mathcal{F}}^{N}$-predictable process such that $\mathbb{P}(\mathrm{d}\omega)$ a.s., $\gamma\in L^{2}([0,T])$. There exists a sequence of random variables:
$\Gamma_{k}:\Omega\rightarrow L^{2}([0,T])$
such that $\Gamma_{k}$ is ${\mathcal{F}}^{N}_{\tau_{k}}$-measurable and
$\mathbb{P}(\mathrm{d}\omega)\times\mathrm{d}t~{}a.s.,\quad\gamma_{t}(\omega)=\sum_{k\geq 0}\Gamma_{k}(t)\mathbbm{1}_{(\tau_{k},\tau_{k+1}]}(t).$
@@
See [5, A2, Th. 34]. ∎
Consider $\mathbb{D}([0,T])$, the Skorokhod space of càdlàg functions on $[0,T]$. Since each random variable $\Gamma_{k}$ is ${\mathcal{F}}^{N}_{\tau_{k}}$-measurable, there exists a deterministic and measurable function $\Psi_{k}:\mathbb{D}([0,T])\rightarrow L^{2}([0,T])$ such that:
$\Gamma_{k}=\Psi_{k}(N_{\cdot\wedge\tau_{k}}).$
This establishes a one-to-one correspondence between an admissible control $\gamma$ and a collection of measurable functions $(\Psi_{k})_{k\geq 0}$. Write =$\mathbb{N}=\{0,1,\cdots\}$ for the set of natural integers. We denote by $\tilde{A}$ the set of all the measurable functions from $D([0,T])$ to $(L^{2}([0,T]))^{\mathbb{N}}$.
In addition, there exists a deterministic function $V:D([0,T])\rightarrow\mathbb{R}$ such that
@@
We will see in Corollary 2.9 below that this function $V$ is explicit.
Given $\Psi\in\tilde{A}$, consider $(\Lambda,N_{t},V_{t})$ which satisfies the three properties outlined in Definition 2.2, with the control given by $\gamma=\sum_{k\geq 0}\Psi_{k}(N_{\cdot\wedge\tau_{k}})\mathbbm{1}_{(\tau_{k},\tau_{k+1}]}$. These three conditions fully determine the law of $(N_{t})_{t\in[0,T]}$. We denote by $Q_{\Psi,\mu}$ this probability law:
$Q_{\Psi,\mu}={\mathcal{L}}((N_{t})_{t\in[0,T]}).$
Thus, our optimization problem (2.1) is equal to:
$\inf_{\Psi\in\tilde{A}}\mathbb{E}\left[\sum_{k\geq 0}\frac{1}{2}\int_{0}^{T}{\Psi^{2}_{k}(N_{\cdot\wedge\tau_{k}})\mathbbm{1}_{(\tau_{k},\tau_{k+1}]}(s)\mathrm{d}s}+\kappa V(N))\right]$
$=\inf_{\Psi\in\tilde{A}}\int\left[\sum_{k\geq 0}\frac{1}{2}\int_{0}^{T}{\Psi^{2}_{k}(\tilde{n}_{\cdot\wedge\tilde{\tau}_{k}})\mathbbm{1}_{(\tilde{\tau}_{k},\tilde{\tau}_{k+1}]}(s)\mathrm{d}s}+\kappa V(\tilde{n}))\right]Q_{\Psi,\mu}(\mathrm{d}\tilde{n}).$
In the last expression, the $\tilde{\tau}_{k}$ denotes the jumps of $\tilde{n}$. The key observation here is that this expression is entirely independent of the original probability space from which we began.

### v2

@@ Abstract
We investigate an optimal control problem motivated by neuroscience, where the dynamics is driven by a Poisson process with a controlled stochastic intensity and an unknown parameter. Given a prior distribution for the unknown parameter, we describe its evolution using Bayes’ rule. We reformulate the optimization problem by applying Girsanov’s theorem and establish a dynamic programming principle. Finally, we characterize the value function as the unique viscosity solution to a finite-dimensional Hamilton-Jacobi-Bellman equation, which can be solved numerically.
@@ 1 Introduction
Let $(Y_{t})$ represent the membrane potential of a spiking neuron. The neuron emits “spikes” randomly at rate $f_{\Lambda}(Y_{t})$. When a spike occurs, the potential resets to zero. Between spikes, the dynamics are governed by the ordinary differential equation
@@
In this model, the functions $(\lambda,y)\mapsto f_{\lambda}(y)\in\mathbb{R}_{+}$ and $y\mapsto b(y)\in\mathbb{R}$ are deterministic and known. The parameter $\Lambda$ is a random variable with a known distribution $\mu$, supported on $\mathbb{R}_{+}$, but its realization is unknown. Our objective is to estimate the value of $\Lambda$ as accurately as possible. The input current $(\gamma_{t})$ is serves as a control and can be chosen to improve the estimation. Crucially, we only observe the spike times of $(Y_{t})$, and thus $(\gamma_{t})$ must depend causally on the spike times up to time $t$. The central question addressed in this article is:
The question we answer in this article can be summarized as follows:
How should the stochastic control $(\gamma_{t})$ be chosen to optimally estimate the unknown parameter?
This leads us to formulate an optimal control problem driven by a counting process with unknown intensity. Our approach extends beyond neuroscience to applications requiring optimal control of processes with unknown parameters, such as in online learning scenarios. We present a mathematical framework to address this class of problems.
Such problems have been extensively studied in the discrete-time stochastic optimal control literature (see, e.g., [8, 10]). A key approach involves assigning a prior distribution to the unknown parameter, which is updated at each time step using Bayes’ rule. This Bayesian framework is central to filtering theory; see [7] for a historical account of filtering theory and [9] for an overview of classical filtering techniques. In this work, we combine insights from filtering theory and stochastic control.
In the continuous-time setting, particularly within the impulse control framework, these issues have been addressed by [2] in a Brownian framework and extended by [1] to a Poisson framework. As in the discrete-time setting, a Bayesian approach is employed: the unknown parameter is sampled from a prior distribution, and a stochastic process in the space of probability measures, known as the posterior distribution, is introduced. This approach leads to a dynamic programming principle and a characterization of the value function via a quasi-variational parabolic equation, interpreted in the sense of viscosity solutions.
In [12] (lectures given at Collège de France) and [4] (with applications to asset management), the authors consider instances of the following diffusive model:
@@
In this model, the process $(X_{t})$ is observed, but the parameter $\Lambda$ is unknown. A key idea introduced in these works is to reformulate the problem using Girsanov’s theorem. This approach eliminates the drift and explicitly accounts for the dependence on the unknown parameter through the exponential martingale in Girsanov’s transformation. Furthermore, in cases where $\Lambda$ appears linearly or for specific classes of prior distributions known as conjugate families, the authors derive a finite-dimensional Hamilton-Jacobi-Bellman (HJB) equation satisfied by the value function.
@@
In the setting we introduce, a novel mathematical challenge arises because the conditional distribution of the unknown parameter evolves continuously and exhibits jumps: these jumps in the conditional distribution occur precisely at the spike times of the underlying point process, with an intensity that depends on the unknown parameter. Another significant challenge arises: modifying the control $(\gamma_{t})$ alters the spike times of the process. Since the controller observes only the spike times, it is not immediately clear which controls are admissible. In particular, it is generally not possible to use the same control $(\gamma_{t})$ for two different initial conditions $y$ and $y^{\prime}$ of $(Y_{t})$. Indeed, changing the initial condition may alter the spike times, rendering a control valid for $y$ invalid for $y^{\prime}$. This feature makes it particularly challenging to obtain a priori regularity estimates for the value function. As in [12, 4], Girsanov’s theorem plays a crucial role in addressing these difficulties.
Another challenge stems from the infinite-dimensional nature of the measure space, where the posterior distribution evolves continuously over time. In cases where the intensity of the underlying point process depends linearly on the unknown parameter, we demonstrate that the posterior distribution is confined to a two-dimensional subspace of $\mathbb{R}^{2}$. Furthermore, we establish that the value function associated with the optimal control problem is the unique viscosity solution to a specific finite-dimensional Hamilton-Jacobi-Bellman (HJB) equation.
We now briefly outline our main contributions and the organization of this paper.
@@
The control problem is introduced in Section 2.1, where we characterize the admissible controls and demonstrate that the optimization problem is independent of the initial probability space. In Section 2.3, we reformulate the problem using Girsanov’s theorem for point processes with stochastic intensity and establish the equivalence between the original and reformulated problems. In Section 2.4, we introduce the value function, and in Section 2.5, we demonstrate how to obtain a finite-dimensional representation of the value function when the function $f_{\lambda}(y)$ is linear in $\lambda$, i.e. $f_{\lambda}(y)=\lambda g(y)$. This result is analogous to the finite-dimensional reduction obtained in [12, 4] for the diffusive setting with linear drift. Thereafter, we focus our analysis on this finite-dimensional problem.
@@
In Section 3.1, we prove that the value function is Lipschitz continuous with respect to the spatial variables. To this end, we conduct a detailed analysis of the exponential martingale involved in Girsanov’s transformation. They key difficulty is that the admissible controls are not assume to be bounded. Using these regularity estimates, we establish a dynamic programming principle with deterministic stopping times in Section 3.2. Leveraging this result, we demonstrate in Section 3.3 that the value function is continuous with respect to time. In Section 3.4, we provide a general dynamic programming principle that accommodates arbitrary stopping times.
@@
In Section 4, we show that the value function is the unique viscosity solution to a specific Hamilton-Jacobi-Bellman (HJB) equation, given by equation (4.1). This first-order HJB equation includes zero-order terms arising from the jumps in the underlying dynamics. We adopt the methodology of [13] to address these terms. A particular challenge in proving the comparison result arises due to the unbounded nature of both the controls and the spatial variables.
@@
Finally, in Section 5, we present several numerical illustrations and discuss the behavior of the optimal control.
@@ 2.1 The optimization problem
Let $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$ be a probability measure supported on $\mathbb{R}_{+}$ with a finite second moment. Consider functions $b$ and $f$ satisfying the following assumption:
@@
The function $b:\mathbb{R}\rightarrow\mathbb{R}$ is Lipschitz, and $f:\mathbb{R}_{+}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ is continuous and bounded for each $\lambda\geq 0$, with quadratic growth in $\lambda$:
$\exists C>0,\quad\sup_{y\in\mathbb{R}}f_{\lambda}(y)<C(1+\lambda^{2}).$
Let $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t}),\mathbb{P})$ be a filtered probability space satisfying the usual conditions. Let $\Pi(\mathrm{d}u,\mathrm{d}z)$ be an $({\mathcal{F}}_{t})$-Poisson measure with intensity given by the Lebesgue measure $\mathrm{d}u\mathrm{d}z$ on $\mathbb{R}_{+}\times\mathbb{R}_{+}$. Let $\Lambda$ be an ${\mathcal{F}}_{0}$-measurable random variable with probability distribution $\mu$. Fix a time horizon $T>0$. Let $(\gamma_{t})_{t\in[0,T]}$ be an $({\mathcal{F}}_{t})$-progressively measurable process such that
@@
We consider the following stochastic differential equation for all $t\in[0,T]$:
$Y_{t}$ $=y+\int_{0}^{t}(b(Y_{u})+\gamma_{u})\mathrm{d}u-\int_{0}^{t}\int_{\mathbb{R}_{+}}Y_{u-}\mathbbm{1}_{\{z\leq f_{\Lambda}(Y_{u}-)\}}\Pi(\mathrm{d}u,\mathrm{d}z).$
This stochastic differential equation, driven by the Poisson measure $\Pi$, admits a unique pathwise solution. Indeed, it suffices to solve the equation between consecutive jumps; see [11, Ch. IV, 9]. We define:
$N_{t}=\int_{0}^{t}\int_{\mathbb{R}_{+}}\mathbbm{1}_{\{z\leq f_{\Lambda}(Y_{u-})\}}\Pi(\mathrm{d}u,\mathrm{d}z),$ (2.1)
as the number of events between 00 and $t$, and denote its canonical filtration by:
${\mathcal{F}}^{N}_{t}=\sigma\{N_{s},~{}s\leq t\}.$
@@
A control $\gamma$ is admissible if it is predictable with respect to the filtration $({\mathcal{F}}^{N}_{t})_{t\geq 0}$.
We denote by ${\mathcal{A}}$ the set of admissible controls. The posterior distribution of $\Lambda$ given the observed spike times is defined as:
$M^{\gamma}_{t}:={\mathcal{L}}(\Lambda~{}|~{}{\mathcal{F}}^{N}_{t}).$
Let $\kappa>0$. The optimization problem is formulated as follows:
In other words, our goal is to minimize the variance of the posterior distribution, with the quadratic cost on the control serving as a regularization to prevent excessively large controls. A key difficulty in this formulation is that the set of admissible controls is not known a priori. Indeed, we require the control $\gamma$ to be predictable with respect to $({\mathcal{F}}^{N}_{t})$, where $N_{t}$ is itself determined by solving a stochastic differential equation driven by a Poisson measure and $\gamma$. In the next section, we provide a more explicit characterization of the admissible controls.
2.2 Admissible controls and structure of the optimization problem
We now reformulate the optimization problem to demonstrate that the value function defined by (2.2) is independent of the probability space $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t}),\mathbb{P})$. Let $(\tau_{k})_{k\geq 1}$ denote the successive jump times of $(N_{t})_{t\geq 0}$, with the convention that $\tau_{0}=0$. The following proposition characterizes the admissible controls.
Proposition 2.3.
Let $\gamma$ be an ${\mathcal{F}}^{N}$-predictable process such that $\mathbb{P}(\mathrm{d}\omega)$ a.s., $\gamma\in L^{2}([0,T])$. There exists a sequence of random variables:
$\Gamma_{k}:\Omega\rightarrow L^{2}([0,T])$
such that $\Gamma_{k}$ is ${\mathcal{F}}^{N}_{\tau_{k}}$-measurable and
$\mathbb{P}(\mathrm{d}\omega)\times\mathrm{d}t~{}a.s.,\quad\gamma_{t}=\sum_{k\geq 0}\Gamma_{k}(t)\mathbbm{1}_{(\tau_{k},\tau_{k+1}]}(t).$
Proof.
See [6, A2, Th. 34]. ∎
Consider $D([0,T])$, the Skorokhod space of càdlàg functions on $[0,T]$. Since each random variable $\Gamma_{k}$ is ${\mathcal{F}}^{N}_{\tau_{k}}$-measurable, there exists a deterministic and measurable function $\Psi_{k}:D([0,T])\rightarrow L^{2}([0,T])$ such that:
$\Gamma_{k}=\Psi_{k}(N_{\cdot\wedge\tau_{k}}).$
Here, we use the notation $a\wedge b:=\min(a,b)$, so that for all $t\in[0,T]$, $N_{t\wedge\tau_{k}}=N_{\min(t,\tau_{k})}$. This establishes a one-to-one correspondence between an admissible control $\gamma$ and a collection of measurable functions $(\Psi_{k})_{k\geq 0}$. We denote by $\tilde{A}$ the set of all the measurable functions from $D([0,T])$ to $(L^{2}([0,T]))^{\mathbb{N}}$.
Proposition 2.4.
Let $\Psi\in\tilde{A}$. Let $\gamma$ be defined by:
$\gamma_{t}=\sum_{k\geq 0}\Psi_{k}(N_{\cdot\wedge\tau_{k}})(t)\mathbbm{1}_{(\tau_{k},\tau_{k+1}]}(t).$ (2.3)
The triple $(\Lambda,(N_{t}),(Y_{t}))$ satisfies the following:
@@ 2.2 Admissible controls and structure of the optimization problem
These properties, together with the relation (2.3), characterize the law of $(N_{t})_{t\in[0,T]}$. We denote this law by $Q_{\Psi,\mu}$:
$Q_{\Psi,\mu}={\mathcal{L}}((N_{t})_{t\in[0,T]}).$
@@
By applying [6, Ch. II, Th. 9], the second established by showing that $N_{t}-\int_{0}^{t}f_{\Lambda}(Y_{s-})\mathrm{d}s$ is an $({\mathcal{F}}_{t})$-martingale. This follows directly from the representation (2.1), using the properties of Poisson measures and the following estimate:
$\mathbb{E}\left[N_{t}\right]=\int_{0}^{t}\mathbb{E}\left[f_{\Lambda}(Y_{s})\right]\mathrm{d}s\leq CT\left(1+\mathbb{E}\left[\Lambda^{2}\right]\right)<\infty.$
Furthermore, the three properties above, combined with the control explicitly defined by (2.3), ensure that $(N_{t},Y_{t})$ can be constructed jump by jump, which guarantees the uniqueness of the law of $(N_{t},Y_{t})_{t\in[0,T]}$. ∎
Moreover, since $\text{Var}(M^{\gamma}_{T})=\mathbb{E}\left[\Lambda^{2}~{}|~{}{\mathcal{F}}^{N}_{T}\right]-\mathbb{E}^{2}\left[\Lambda~{}|~{}{\mathcal{F}}^{N}_{T}\right]$ is ${\mathcal{F}}^{N}_{T}$-measurable, there exists a deterministic function $V:D([0,T])\rightarrow\mathbb{R}$ such that:
@@
In Corollary 2.9 below, we show that the function $V$ admits an explicit expression.
Thus, our optimization problem (2.2) can be expressed as::
$\inf_{\Psi\in\tilde{A}}\mathbb{E}\left[\sum_{k\geq 0}\frac{1}{2}\int_{0}^{T}{\Psi^{2}_{k}(N_{\cdot\wedge\tau_{k}})(s)\mathbbm{1}_{(\tau_{k},\tau_{k+1}]}(s)\mathrm{d}s}+\kappa V(N)\right]$
$=\inf_{\Psi\in\tilde{A}}\int\left[\sum_{k\geq 0}\frac{1}{2}\int_{0}^{T}{\Psi^{2}_{k}(\tilde{n}_{\cdot\wedge\tilde{\tau}_{k}})(s)\mathbbm{1}_{(\tilde{\tau}_{k},\tilde{\tau}_{k+1}]}(s)\mathrm{d}s}+\kappa V(\tilde{n})\right]Q_{\Psi,\mu}(\mathrm{d}\tilde{n}).$
Here, $\tilde{\tau}_{k}$ denotes the jump times of $\tilde{n}$. The key observation is that this expression is entirely independent of the original probability space.